OpenStudy (anonymous):

What is the solution to the system? (Use elimination is possible) -18.6x+17.2y=90.2 14.2x+3.4y=35.8

OpenStudy (the_fizicx99):

Okie, -18.6x+17.2y=90.2 14.2x+3.4y=35.8 You want to cancel out a variable to solve for then to solve for the other. So, basically multiply every number by 14.2 so, -18.6x(14.2)+17.2y(14.2)=90.2(14.2) and 14.2x(18.6) + 3.4y(18.6) = 35.8(18.6) So over all its, -264.12x + 244.24y = 1,280.84 264.12x + 63.24y = 665.88 Combine these two together, the x's cancel out so your left with: 307.48y = 1,946.72 Divide both sides by 307.48 to isolate y and then plug in what you got for y to solve for x.